### v1

# Tunable quasi-discrete spectrum of spin waves excited by periodic laser patterns

Abstract
We present a concept for selective excitation of magnetostatic surface waves with quasi-discrete spectrum using spatially patterned femtosecond laser pulses inducing either ultrafast change of magnetic anisotropy or inverse Faraday effect. We micromagnetically simulate excitation of the waves with periodically patterned uni- or bipolar laser impact. Such excitation yields multiple wavepackets propagating with different group velocities, whose dispersion corresponds to the set of quasi-discrete points. In addition, we show that the frequency of the spectral peaks can be controlled by the polarity of the periodic impact and its spatial period. Presented consideration of multiple spatially periodic SW sources as a whole enables implementation of a comprehensive toolkit of spatio-temporal optical methods for tunable SW excitation and control of its parameters.
Spin waves, magnonics, micromagnetics, ultrafast magnetism
I Introduction
Recent progress in magnonics demonstrates multiple ways to use spin waves (SW) as data carriers for non-charge-based energy-efficient computation Flebus et al. (2024). Implementing SW offers wide possibilities for processing information, when conventional Boolean operations are realized in magnonic integrated circuits Mahmoud et al. (2022); Wang et al. (2020); Garlando et al. (2023). This promotes THz operation rates as well as scaling down the devices to the nanometer range comparable to the wavelengths of SW Wang et al. (2023); Hortensius et al. (2021). Furthermore, the lack of Ohmic losses during SW propagation, the wide available material range, and the potential to employ non-boolean wave-based calculations make magnonics a highly promising base for the new-generation computing Mahmoud et al. (2020); Papp, Porod, and Csaba (2021); Yaremkevich et al. (2023).
The use of coplanar antennas for excitation and detection of SW as well and its combination with Brillouin light scattering is now considered as a standard in magnonics. This approach revealed a variety of effects of SW interaction to be implemented in logic devices, when SW is excited continuously Talmelli et al. (2020); Wang et al. (2020); Sadovnikov et al. (2015, 2022) or by nanosecond microwave pulses Breitbach et al. (2024); Mahmoud et al. (2022). However, achieving of the highest performance requires an excitation of SW wavepackets of sufficiently short duration, which may increase the clock speed of information processing. Femtosecond laser pulses have been utilized for controllable excitation of such SW wavepackets since the first demonstrations Satoh et al. (2012); Van Kampen et al. (2002); Au et al. (2013a); Yoshimine et al. (2014). Further exploration of various means of active control of parameters of laser-induced SW, such as phase and group velocities, is required for developing optomagnonic concepts Bublikov et al. (2023); Kolosvetov et al. (2022).
Optically reconfigurable magnonics uses optically-induced changes of material parameters to tune SW properties Vogel et al. (2015). Combined with femtosecond laser pulses, it offers comprehensive tools for SW investigation in the spatial, temporal, and frequency domains with sub-micrometer and sub-picosecond resolutions Chernov et al. (2020); Hortensius et al. (2021); Hashimoto et al. (2017); Khokhlov et al. (2019). Moreover, optical methods allow one to shape excitation in a spatially non-trivial form, and thus tune wavenumber distribution of SW Satoh et al. (2012). Recently, the interaction of laser-induced SW from a pair of localized sources was demonstrated Yoshimine et al. (2017); Kolosvetov et al. (2022); Kolosvetov and Chernov (2023). The development of a model describing composite spatially patterned laser excitation as a whole and the analysis of associated magnetization dynamics both inside and outside the excitation pattern can simplify its practical implementation and can reveal additional ways to control SW parameters Talmelli et al. (2020).
In this article, we develop the concept of selective excitation of magnetostatic surface waves (MSSW) with quasi-discrete spectrum using spatially patterned femtosecond laser pulses. The proposed spatial distributions of the intensity and polarization of excitation pulses can be obtained experimentally using a spatial light modulator, as used, for example, in Refs. Kolosvetov et al. (2022); Kolosvetov and Chernov (2023). We consider a model of MSSW excitation due to the two experimentally verified mechanisms: ultrafast thermal change of magnetic anisotropy (CMA) and inverse Faraday effect (IFE), and we validate this model in the case of single pulse excitation using micromagetic simulations. Analyzing the dispersion of thus excited MSSW, we show no significant difference between waves propagating beyond the excitation area. Second, we micromagnetically simulate excitation of MSSW with periodically patterned laser pulses. Periodic excitation via IFE is considered to be either uni- or bipolar, depending on the modulation of the helicity of the excitation pattern, while CMA excitation is unipolar only. Considering pump pulses periodically distributed in space as a single SW source, we show that it has a quasi-discrete periodic spectrum in reciprocal space with either even or odd modes. Indeed, MSSW characteristics outside the excitation area are governed by the polarity and periodicity of the excitation pattern. Inside the pattern, the excitation mechanism dictates the spin wave evolution. Thus, upon long-lived CMA effect, periodicity of the pump in space leads to the formation of a transient reconfigurable magnonic crystal structure Vogel et al. (2015).
II Model of MSSW excited by periodic laser pattern
II.1 Geometry of MSSW laser excitation
To construct the model of spatially periodic laser excitation of MSSW, we consider bismuth-doped yttrium iron garnet Bi:YIG as a magnetic material, since it supports CMA and IFE for excitation of magnetization dynamics Shelukhin et al. (2018). As we are interested only in the excitation of MSSW, we reduce the model to one-dimensional (1D) geometry (Fig. 1 a), and consider a film, where plane MSSW are excited by a laser spot limited in the $x$ direction and infinitely long in the $y$ direction, as in experiments with elongated pump laser spots Satoh et al. (2012); Kainuma, Matsumoto, and Satoh (2021); Matsumoto et al. (2020); Hioki, Hashimoto, and Saitoh (2020, 2022); Yoshimine et al. (2017). In-plane external magnetic field $\textbf{H}_{\text{ext}}$ is oriented along the $y$ axis, and MSSW propagate perpendicularly to $\textbf{H}_{\text{ext}}$, i.e. along $x$ axis. In our model, the uniaxial magnetic anisotropy axis with the equilibrium parameter $K_{u}$ is oriented along $x$ axis to fulfill the condition of optical excitation of MSSW due to CMA Khokhlov et al. (2019); Filatov et al. (2020). Also, we assume the excitation to be uniform across thickness of the film to exclude the exchange standing spin waves (SSW).
To describe periodically patterned excitation of MSSW due to either CMA and IFE, we introduce the spatio-temporal profile of laser impact as:
$I(x,t)=I_{0}G(x)T(t),$ (1)
where $I_{0}$ corresponds to the amplitude of the impact; $G(x)$ and $T(t)$ describe spatial and temporal profiles of the impact, respectively. We assume that the individual pump spot has a Gaussian profile in space. Thus, the total spatially periodic excitation profile is determined by 1D function of $12N+$ 1 Gaussians centered at $x=0$ and periodically spaced by a distance $\delta$ (upper panel in Fig. 1 b):
$G_{\text{uni}}(x,N,\delta)=\sum_{n=-N}^{N}\exp\left[-\frac{(x+n\delta)^{2}}{2\sigma^{2}}\right],$ (2)
where $\sigma=\text{FWHM}/(2\sqrt{2\ln{2}})$ ), and full width at half maximum (FWHM) is fixed to 1 $\mu m$ being larger than the diffraction limit to fulfill experimental conditions Khokhlov et al. (2019); Filatov et al. (2022); Iihama et al. (2016). Equation (2) describes a single Gaussian function if $N=0$, and adds ±$\pm N$ spatially periodic Gaussian functions located on both sides of the central one with separation by a period $\delta$. Such unipolar pattern is applicable for both CMA excitation and IFE excitation with a fixed circular polarization of laser pulses To account for the bipolarity of IFE when the polarization of the laser pulses is altered between the neighbouring spots Kimel et al. (2005), Eq. (2) is modified as (upper panel in Fig. 1 c):
$G_{\text{bi}}(x,N,\delta)=\sum_{n=-N}^{N}(-1)^{n}\exp\left[-\frac{(x+n\delta)^{2}}{2\sigma^{2}}\right].$ (3)
The temporal profile $T(t)$ of the impact Eq. (1) depends on the physical process induced by the laser in the material, as discussed below.
Figure 1: The geometry of the plane MSSW excitation with periodic pump pattern and the corresponding spectral distribution. a, the geometry under investigation. b, c, the distributions of the Fourier amplitudes calculated analytically using (6) for uni- and bipolar impact, respectively. The spatial profiles are illustrated above the corresponding maps. The numbers in white denote the even and odd order of the modes with corresponding values of wavenumber $k$ indicated by solid red lines. Horizontal black dashed lines illustrate the values $\sigma$ and $3\sigma$. The horizontal blue dashed line corresponds to $\delta=3$$\mu$m used in the following simulations.
II.2 Physical impacts of the laser pulse
Laser-induced CMA is introduced as an ultrafast reduction of the equilibrium uniaxial magnetic anisotropy parameter $K_{u}$ with the maximal relative amplitude $\Delta K_{u}$. In dielectric magnets, the CMA occurs on the picosecond time scale Shelukhin et al. (2018), which is 2-3 orders of magnitude faster than the MSSW period Chernov et al. (2017). In turn, relaxation processes are on the nanosecond time scale, which is comparable to the typical propagation time of magnetostatic waves Satoh et al. (2012); Matsumoto et al. (2020); Kainuma, Matsumoto, and Satoh (2021). Thus, we approximated the temporal dependence of laser-induced CMA by the Heaviside step function $T_{H}(t)=\Theta(t-t_{0})$, where $t_{0}$ is the moment of pump impact. The resulting spatio-temporal profile of periodically patterned CMA is:
$K_{u}^{\prime}(x,t,N,\delta)=K_{u}\left[1-\Delta K_{u}G_{\text{uni}}(x,N,\delta)T_{H}(t)\right].$ (4)
The laser-induced IFE in our model is introduced as a pulse of the effective out-of-plane field $\textbf{H}_{\text{IFE}}$, according to experimental observations Kimel et al. (2005); Kirilyuk, Kimel, and Rasing (2010). The field follows the temporal envelope of femtosecond laser pulse and is described by the Gaussian profile $T_{G}(t)=\exp\left[-(t-t_{0})^{2}/\left(2\sigma_{t}^{2}\right)\right]$, where $\sigma_{t}=\text{FWHM}_{t}/(2\sqrt{2\ln{2}})$ ) and FWHMt is set to 200 fs to meet the typical experimental pump pulse duration Yaremkevich et al. (2021); Khokhlov et al. (2019); Hioki, Hashimoto, and Saitoh (2020); Krichevsky et al. (2024). In addition, we take into account the possibility to control the polarity of $\textbf{H}_{\text{IFE}}$ by the choice of right- or left-handed circular polarization of pump pulses in experiments Kimel et al. (2005); Kirilyuk, Kimel, and Rasing (2010). Thus, the periodically patterned IFE in our model is represented by unipolar (2) or bipolar (3) impact in space. The resulting spatio-temporal profile of the IFE with maximal amplitude $H_{\text{IFE}}$ takes the form:
$H_{\text{IFE}}^{\prime}(x,t,N,\delta)=H_{\text{IFE}}G(x,N,\delta)T_{G}(t),$ (5)
where $G(x,N,\delta)$ has the form (2) or (3).
Note that the approximation  (4) works for thermally induced CMA, which only has the unipolar impact (2). However, there are other mechanisms of CMA in iron garnets, including non-thermal ones Kirilyuk, Kimel, and Rasing (2010); Stupakiewicz et al. (2017) with alternating sign. Laser-induced demagnetization of a ferromagnet usually appears along with CMA Gerevenkov et al. (2021); Maehrlein et al. (2018). However, it has much longer rise times in iron garnets and has a value of an order of magnitude smaller than CMA Shelukhin et al. (2018); Davies et al. (2019); Deb et al. (2021). Thus, we omitted the demagnetization process in the model.
The approximation of the excitation being homogeneous across the thickness is valid for iron garnets with thicknesses ranging from tenths to hundreds of nanometers. In general, however, pump pulse impact in iron garnet films is non-uniform across the thickness in accordance with light penetration depth. This should cause the excitation of SSW or higher-order modes of MSSW. We omit this from the consideration in order to focus on propagation of lowest-order MSSW. We note that high-quality ultrathin garnet films where our assumption is valid are now becoming available Levy et al. (2019); Hartmann et al. (2024).
II.3 Analysis of wavenumber distribution
To analyze the excitation of the MSSW via CMA and IFE, we investigate the distribution of the wavenumbers $k$, excited by such a periodically patterned impacts. The wavenumber distribution of the laser-induced SW amplitudes is known to be determined by the Fourier transform of the pump spatial profile Satoh et al. (2012); Hortensius et al. (2021). We find analytically the Fourier amplitudes as a function of wavenumber for uni- and bipolar spatially periodic patterns (2,3). Then, the Fourier transform for arbitrary $N$ has the form:
$\widetilde{A}(k,N,\delta)=\sqrt{\frac{\cos{\left[k\delta(2N+1)\right]}\mp 1}{\cos{\left(k\delta\right)}\mp 1}}\sigma\exp\left(-\frac{k^{2}\sigma^{2}}{2}\right),$ (6)
where top and bottom signs correspond to the uni- and bi-polar excitation, respectively. To demonstrate the main findings, we limited the total number of Gaussian functions to 5 ($N=2$) in further consideration. In Eq. (6), the first factor is a Dirac comb Kindermann and Mele (2011) for finite $N$. The spatially limited Dirac comb we propose also demonstrates the same periodicity, but allows us to consider wave propagation both inside the pattern and outside it.
We calculate the corresponding Fouirer amplitude distribution of the modes as the dependence on the spatial period of the pump pattern $\delta$ (Fig. 1 b and c). In the case of unipolar impact, the distribution has maximal amplitudes around even order modes starting at $k=0$ (Fig. 1 b). In contrast, the bipolar impact leads to emergence of odd order modes (Fig. 1 c). This distribution is multiplied by the overall decrease in amplitude with $k$ governed by the spatial width $\sigma$ of the individual laser spot in the pattern and described by the exponential factor in Eq. (6). The spacing between the spots $\delta$ controls both the wavevnumbers and the amplitudes of the excited modes, as seen in Fig. 1 b and c. For example, increasing $\delta$ leads to excitation of higher-order modes, as their wavenumbers decrease and they fall within a range of $k$ limited by the size of the single spot. Thus, while the total width of the excited wavenumber is determined by $\sigma$, the periodicity establishes the quasi-discrete set of wavenumbers with spectral positions controlled by the spatial period $\delta$. As expected, the quasi-discrete spectrum emerges if $\delta$ exceeds $3\sigma$, and the excitation pattern acquires a contrast higher than $30\%$. We also note that the change of the total number of Gaussians in periodic pattern affects their spectral width (see Appendix B).
The established model outlines a possibility to control SW parameters using periodically patterned laser excitation in addition to the tunability of SW laser excitation based on pump spot shape and size Satoh et al. (2012); Chernov et al. (2017), mutual orientation of the external magnetic field and magnetic anisotropy axes Khokhlov et al. (2019); Khokhlov, Filatov, and Kalashnikova (2024). Additional tools are a spatial period of the pattern $\delta$, a choice of excitation mechanism, CMA or IFE, as well as a choice between uni- and bipolar excitation as seen from Eq.(6) and Fig. 1 b and c. According to our model, CMA being always unipolar impact offers an additional way to control SW parameters. In particular, CMA is a long-living effect compared to MSSW propagation, which results in the formation of periodic non-uniformity of $K_{u}$ in space. Thus, one expects that the part of excited MSSW can stay inside this structure, filling it as a magnon resonator.Further, we support the above analytical predictions with the results of micromagnetic modeling.
Figure 2: MSSW excitation by the single spatial Gaussian impact. The panels a-d and e-h correspond to the excitation by the CMA and IFE, respectively. a, e, Spatial profiles of the excitation mechanisms (left), and the spatio-temporal distribution of the out-of-plane magnetization component $M_{z}(t,x)$ in the vicinity of the excitation area (right). b, f, $M_{z}(t,x)$ dynamics throughout the entire range of time and distance. c, g, Frequency spectra of magnetization dynamics at various distances from the excitation area obtained by 1D FFT of $M_{z}(t)$ at fixed coordinates $x=$ 0, 10, 20 $\mu$m depicted with dashed lines of corresponding color in panels b and f. The black dashed lines show the high-frequency maximum.d, h, Excited wave dispersions obtained by 2D FFT of the $M_{z}(t,x)$ data in panels b and f.
III Micromagnetic simulations
III.1 Modeling parameters
Micromagnetic validation of the proposed model is performed using Ubermag Beg, Lang, and Fangohr (2022) with object oriented micromagnetic framework (OOMMF) Donahue and Porter (1999) as a micromagnetic calculator. In simulations, we use the material parameters, typical for Bi:YIG films Zeuschner et al. (2022): magnetization saturation $\mu_{0}M_{s}=175$ mT, exchange energy constant $A_{ex}=3.5\times 10^{-12}$ J/m, Gilbert damping $\alpha=0.01$, and uniaxial anisotropy parameter $K_{u}=5750$ J/m3. The in-plane external magnetic field $\mu_{0}H_{\text{ext}}=100$ mT is applied with a tilt of 1 with respect to $Y$ axis. We set the lateral size of the computational cell as $10\times 10$ nm2 to be smaller than the magnetostatic exchange $\sqrt{2A/(\mu_{0}M_{s}^{2})}=16.9$ nm and the magnetocrystalline exchange lengths $\sqrt{A/K_{u}}=24.7$ nm both Abo et al. (2013). The thickness of the film is 500 nm. The corresponding sagittal cell size is 500 nm to exclude the excitation of SSW and higher-order MSSW. 1D periodic boundary conditions are applied along $Y$ axis. In the simulations, time-domain dynamics is tracked by solving the Landau-Lifshitz-Gilbert equation with the set of energy terms: Zeeman energy, exchange energy, energy of the uniaxial magnetic anisotropy, and magnetostatic energy. CMA is simulated by introducing the spatially non-uniform parameter $\widetilde{K_{u}}(x,t,N,\delta)$ described by (4). IFE is simulated by adding to the system the Zeeman energy term with the transient field in the form (5). The amplitudes of CMA and IFE are close to the experimentally achievable values: $\Delta K_{u}=0.3$ Carpene et al. (2010); Gerevenkov et al. (2021) and $\mu_{0}H_{\text{IFE}}=\pm 1$ T Kimel et al. (2005); Kalashnikova, Kimel, and Pisarev (2015); Kozhaev et al. (2018); Shelukhin et al. (2018). The ground state of the magnetization is calculated as a relaxed state after the initializing of the system with the equilibrium parameters, minimization of its energy, and time domain calculations during 10 ns with an increased Gilbert damping of $\alpha=1$ without excitation.
III.2 Single pulse excitation
First, we consider the excitation of MSSW using a single source, i.e., at $N=0$ in Eq. (2). In the spatio-temporal domain, it is common to monitor the out-of-plane magnetization component $M_{z}$, following the optical pump-probe experiments using Faraday Satoh et al. (2012); Chernov et al. (2017) or polar Kerr Khokhlov et al. (2019); Filatov et al. (2022); Au et al. (2013b) effects. CMA and IFE are used as excitation mechanisms, and the corresponding spatial excitation profiles at $t=t_{0}$ are shown in the left panels of Fig. 2 a and e. The right panels show the corresponding spatio-temporal dependencies of $M_{z}(t,x)$. Within the excited area ($|x|<\sigma$), the laser-induced magnetization precession is observed at $t>t_{0}=25$ ps. The propagation of the MSSW is observed away from the excitation area in the form of a wavepacket. Regardless of the excitation mechanism, the wave phase and group velocities of the same sign are characteristic for MSSW, and waves propagate over a distance of more than 25 $\mu$m within the modeled time window of 4.5 ns (Fig. 2 b and f). In addition, the initial phase of the waves is defined by the circular polarization direction in the case of IFE Kimel et al. (2005); Satoh et al. (2012),
However, upon closer examination, the differences in the dynamics of $M_{z}(t,x)$ are observed between the two excitation mechanisms (right panels of Fig. 2 a and e): a sharper phase front evolution around $x=0$ in the IFE case indicates differences in frequencies between the two mechanisms. To explain it, we consider the FFT of $M_{z}(t)$ traces at various distances from the excitation area (Fig. 2 c and g). At $x=0$ (red lines), an excitation of a wide frequency range is observed with two peacks near 2 and 4 GHz for both IFE and CMA. The low-frequency peak is more pronounced in the IFE case, and the high-frequency peak has a higher frequency in the CMA case (marked with a dashed line). Out of the excitation area (blue and green lines in Fig. 2 c and g), the frequency spectra are similar for the two excitation methods and lie between the frequencies of the initial peaks. The wave dispersions for the two excitation mechanisms obtained by 2D FFT (Fig. 2 d and h) indicate that the two peaks correspond to the minimum and maximum frequencies of MSSW.
The observed differences in the excited spectra are due to the different time dependences for the two excitation methods. In the case of IFE (5) the laser-induced field disappears in a time much shorter than the characteristic period of magnetization dynamics. After pumping, the excited dynamics at $x=0$ has a certain frequency spectrum and then MSSW propagates out of the excitation area with group velocities corresponding to its dispersion. In the case of CMA Eq. (4) the material parameter $K_{u}$ experiences a long-lived partial quenching in the pump area. As condition $\mu_{0}H_{\text{ext}}>2K_{u}/M_{s}$ is satisfied, CMA shifts the dispersion curve to higher frequencies within the excitation area Filatov et al. (2022, 2020); Khokhlov et al. (2019). As a result, the state with minimum frequency inside the excited area corresponds to the states with non-zero group velocity outside this area. This leads to a less pronounced low-frequency peak in the excitation area in the case of CMA (red line in Fig. 2 c). In contrast, the high-frequency part experiences a magnon ”trap” due to CMA, since there are no magnon states at these frequencies outside the pumped area. It leads to an increase in the maximum excited frequencies in the case of CMA and a more pronounced component with the wavenumbers $|k|>5$ rad/$\mu$m observed inside the excitation area (Fig. 2 d). Thus, the differences in the spectra are governed by the parts of the dispersion with near-zero group velocity and are the most pronounced near the excitation area (see Appendix A for details). The propagating MSSW outside the pumped area can be considered identical for CMA and IFE.
III.3 Periodically patterned excitation of MSSW
Figure 3: MSSW excitation by spatially periodic Gaussian impacts with $52N+1=$ 5. The rows (a,b), (c,d) and (e,f) correspond to the excitation by CMA, uni- and bipolar IFE, respectively. a, c, e, Spatial profiles of excitation mechanisms (left), and the spatio-temporal distribution of $M_{z}(t,x)$ in the vicinity of the excitation area (right). b, d, f, $M_{z}(t,x)$ dynamics throughout the entire range of time and distance.
Figure 4: Fourier analysis of MSSW excited by spatially periodic Gaussian impacts with $52N+1=$ 5. The rows a-c, d-f and g-i correspond to the excitation by CMA, uni- and bipolar IFE, respectively. a, d, g, Frequency spectra of magnetization dynamics at various distances from the excitation area obtained by 1D FFT of $M_{z}(t)$ at fixed coordinates $x=$ 0, 1.5, 10 $\mu$m. The coordinates are depicted with dashed lines of corresponding color in panels b,e and h. c, f, i, Excited wave dispersions obtained by 2D FFT of $M_{z}(t,x)$ data in Fig. 3 b,d and f, respectively. Horizontal black dashed lines in all panels indicate the same frequencies in the spectra. In the dispersions c, f, i orders of modes are specified and the corresponding wavenumber distributions calculated with (6) are shown by solid blue lines.
[…]
In all three cases, right after the excitation, the phase front is periodic in space and follows the excitation pattern. In the case of IFE, such a spatially periodic phase is also observed at longer time delays because the material preserves its homogeneous magnetic characteristics. The apparent change of the phase front at longer delay times is a result of constructive and destructive interference of the MSSW. When CMA excitation is considered, there is an evident evolution of the phase front already after 1 temporal period. As seen in the right panel of Fig. 3 a, the phase front within the excitation spots and between them becomes inhomogeneous. Furthermore, the temporal period of the oscillations of $M_{z}$ changes with time (see the temporal evolution at $x=0,\pm 3,\pm 6$$\mu$m). This intricate behavior can be understood as an evolution of the spin-wave distribution inside a set of magnon resonators, which eventually results in a establishing of MSSW dispersion characteristic to a magnonic crystal.
The distribution of $M_{z}(t,x)$ throughout the spatio-temporal area demonstrates the propagation of multiple MSSW wavepackets with different group velocities (Fig. 3 b,d and f). The group velocities of the wavepackets are approximately the same when excited by CMA and unipolar IFE (Fig. 3 b and d). The change in IFE polarity results in the excitation of MSSW wavepackets with decreased group velocities (Fig. 3 d and f). Although the temporal period of magnetization dynamics appeared to be the same within the excitation area for the uni- and bipolar IFE, the temporal period of the outgoing wavepackets decreases when the bipolar IFE is simulated.
We performed 1D FFT of $M_{z}$ temporal traces at each $x$ coordinate to show the evolution of the spectra when the MSSW wavepackets propagate outside the excitation area. The FFT spectra inside the central excitation spot at $x=0$ (red lines), between the spots at $x=1.5$$\mu$m (blue lines) and outside the excitation area at $x=10$$\mu$m (green lines) are shown for the CMA, uni- and bipolar IFE excitation in Fig. 4 a,d and g, respectively. These spectra are cross sections of the corresponding spatial evolution of the spectra (Fig. 4 b,e and h) at certain coordinates $x$. Also, in Fig. 4 c,f and i, we show the respective dispersion characteristics of the excited MSSW obtained via 2D FFT of $M_{z}(t,x)$ maps, shown in Fig. 3 b,d and f. Analysis of these spectra allows us to identify the excited MSSW modes and analyze their spectral composition in detail. In particular, all spectra have a quasi-discrete character (Fig.4 c,f and i). From a comparison of the calculated MSSW dispersion with the spectrum obtained from the analytical model Eq. (6), it is seen that the peaks in the spectra correspond to the predicted modes. In the case of CMA and unipolar IFE these are only even modes, while in the case of bipolar IFE – only odd modes. However, details of the spectra differ depending on the excitation mechanism.
[…]
Secondly, the unipolar IFE excitation results in the same set of even-order wavenumbers (Fig. 4 f). The main difference lies in the narrower total width of the dispersion and in the significant decrease of FFT amplitude of the zero-order mode, compared to the higher-order modes at any $x$. The latter, as a result, leads to a decrease in the amplitude of the outgoing wavepacket compared to the second-order mode (Fig. 4 e). The reason of this is explained in terms of CMA and IFE temporal profiles, i.e. by the formation of the magnon trap, as discussed in the previous section. Furthermore, under CMA excitation the low-frequency near-zero-$k$ mode is localized initally inside the resonator between the excitation spots at $x=1.5$$\mu$m (blue line in Fig. 4 a and blue halos slightly above the corresponding horizontal black dashed line around $x=0$ and $k=0$ in Fig. 4 b and c). In turn, under IFE excitation, this mode is nearly absent within at the same position (Fig. 4 d-f). This confirms the possibility of inducing a magnon resonator when the CMA excitation is chosen instead of IFE.
In the case of bipolar IFE excitation, the spectra show the expected odd order MSSW modes (Fig. 4 i). Although low-frequency near-zero-$k$ mode is presented in the spectra at $x=0$, its amplitude is small (Fig. 4 g), and its propagation is not observed. Instead, odd-order modes starting from the 1st propagate with lower group velocities compared to the unipolar IFE case, according to dispersion. Thus, the choice of the IFE polarity allows one to control the order of the excited modes.
IV Conclusion
We presented a model for spin wave excitation, which describes the laser impact on the magnetic system in a single-pulse geometry and extended it to the case of periodically patterned Gaussian laser pulses. The results demonstrate the way to excite multiple MSSW wavepackets propagating with different group velocities, whose dispersion corresponds to the set of selected quasi-discrete points. The choice of the excitation mechanism, CMA or IFE, allows one to set whether the magnon resonator is induced and the localized modes are excited or not. By considering the patterned SW source, we demonstrated that in reciprocal space it corresponds to the excitation of even or odd modes, the frequencies of which can be controlled by the polarity of the periodic impact and its spatial period.
The developed model proves and extends the flexible reconfigurability offered by optical methods of SW excitation. Indeed, tunable periodical patterning of the spatial profile of the laser pulse can be achieved, for example, in the double pump geometry used for the creation of transient gratings Janušonis et al. (2016); Carrara et al. (2022) or by a spatial light modulator Kolosvetov et al. (2022); Kolosvetov and Chernov (2023). The model may find practical implementation in the development and modeling of complicated magnonic systems with multiple SW sources, which is one of the required conditions for the design of magnon logic devices Talmelli et al. (2020). Also, the presented approach may be useful to describe magnonic neuromorphic computing devices, as it offers a consideration of the system with a number of SW sources. We note that our model can be readily improved to include in the consideration 2- and 3-dimensional spatially patterned SW sources and to excite and control other types of SW. As for the next steps, we suggest the temporal modulation of the laser impact in magnonic systems, in addition to the spatial patterning considered here, as it makes possible to implement a comprehensive toolkit of spatio-temporal methods for tunable SW excitation and control of its parameters Jäckl et al. (2017); Muralidhar et al. (2020); Hula et al. (2022).
ACKNOWLEDGMENTS
The work was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics ”BASIS” (grant No 22-1-5-122-1).
COMPETING INTERESTS
The authors declare no competing interests.
author contributions
Iaroslav A. Filatov: Conceptualization, Methodology, Investigation, Software, Writing – Original draft preparation (equal), Writing – Reviewing and Editing (equal). Petr I. Gerevenkov: Data curation, Discussion (equal), Visualization, Writing – Original draft preparation (equal). Nikolai E. Khokhlov: Supervision (equal), Discussion (equal), Writing – Reviewing and Editing (equal). Alexandra M. Kalashnikova: Supervision (equal); Writing – Reviewing and Editing (equal).
Data availability statement
Data supporting the findings of this study are available from the corresponding author upon reasonable request.
Appendix A Spectral difference between CMA and IFE
The Fourier image in the time domain of a femtosecond pulse with $\text{FWHM}_{t}=200$ fs has the form of a Gaussian function with $\sigma_{f}=\left(2\pi\sigma_{t}\right)^{-1}$ and the corresponding $\text{FWHM}_{f}\approx 4.4$ THz. This corresponds to a close to uniform distribution of excited amplitudes over the entire GHz frequency range under consideration. Thus, the frequency spectrum of magnetization dynamics in the excitation area is determined by: (i) the Fourier image of excitation in space (see Sec. II.3) and (ii) the density of states determined by the dispersion of magnetostatic waves (see Figs. 2 (d and h) and 4 (c,f,i)) and inversely proportional to the group velocity. For a medium with selected parameters (see Sec. III.1) in the excited range of wave vectors, the dispersion exhibits a maximum of states near the frequency of 4 GHz, which explains the predominance of this part of the spectrum over the FMR frequency of about 2 GHz [see red lines in Figs. 2 (c and g) and 4 (a,d,g)].
Appendix B Wavenumber distribution dependence on number of Gaussians $N$
Figure 5: Wavenumber distribution at different $N$. a and b, uni- and bipolar cases, respectively.
Increase of $N$ does not lead to a change in the period and positions of the main peaks of the Fourier spectrum, while makes the main peaks narrower (Fig. 5).